rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_blue
icon_green
 

FS#1786 — FS#5763 — Optical switching buckle

Attached to Project— ADSL / SDSL
Modernization
Backend / Core
CLOSED
100%
We are going to launch the first optical buckle tonight around 11h00. that's why we have to cut the NRA, PIL, BLE, BOI, FAC, SEC and LIB.

This launching is expected during an important part of night when internet access are cut.

Thank you for your understanding !

We will keep you informed of all different steps.

Date:  Monday, 12 September 2011, 15:55PM
Reason for closing:  Done
Comment by OVH - Friday, 09 September 2011, 01:43AM

Optical fibers are switched on the optical chassis. We are starting work to mount the optical rings.



Comment by OVH - Friday, 09 September 2011, 01:44AM

PIL and LIB sites are attached optically on the POP of Roubaix.


Comment by OVH - Friday, 09 September 2011, 07:47AM

SEC FAC and BOI sites are related optically. We are setting the BLE site in order to include it in the optical buckle. We are re-establishing meanwhile the traffic.


Comment by OVH - Friday, 09 September 2011, 07:50AM

All optical equipments are implemented. We need a link to finalise the buckle. We faced a problem while implementation. All sites are fixed and calibrated. We will accomplish the routing and work tonight.


Comment by OVH - Friday, 09 September 2011, 07:51AM

NRA, LIB, SEC, FAC, BOI, BLE and PIL Routing are again UP.


Comment by OVH - Friday, 09 September 2011, 18:49PM

There was latency periods and packets losses on the NRA of BLE59.
We intervened on the spot to change certain passive euipments on the site.

We have a better one but we still have some problems.

We are still looking at the level of the routing.


Comment by OVH - Friday, 09 September 2011, 18:52PM

The problem is at the level of the routing.

We are changing the router. The wiring is in progress.
All the NRA will be recables in the following minutes.